Anne and Nancy use a metal alloy that is 21% copper to make jewelry. How many ounces of 19% alloy must be mixed with a 25% alloy to form 78 ounces of the desired alloy?
Answer by wuwei96815(245)

Let x=the 19% alloy and 78-x the 25% alloy. Now, we should be able to write the following equation:

19%x + 25%(78-x) = 21%(78)
.19x + 19.5 - .25x = 16.38
.19x - .25x = 16.38 -19.5
-.06x = -3.12
x = 52
78-(x) = 78-52 = 26

So, you need 52 ounces of the 19% and 26 ounces of the 25%. I hope that is right.